Happy Valley Land and Snow Company provides the following ABC costing information: Activities Total Costs Activity cost drivers Labour hours $160,000 Gas 8,000 hours $18,000 6,000 gallons Invoices20,000 Total costs 198,000 The above activities used by their three departments are as follows 2,500 invoices Bush Dept. 1,200 hours 800 gallons Plowing Dept. 4,300 hours 3,700 gallons Lawn Dept. Labour hours 2,500 hours Gas Invoices How much of the total cost will be assigned to the Lawn Dept? 1,500 gallons 1,600 invoices 400 invoices 500 invoices o a. $29,600 O b. $67,300 c.$100,000 d. $50,000
Answer #1
Lawn Dept.
Labour hours 50000 =160000/8000*2500
Gas 4500 =18000/6000*1500
Invoices 12800 =20000/2500*1600
Total 67300
Option B is correct